OpenStudy (anonymous):

Help! I've been trying to do this for the pas hour or so. I've been using the Pythagorean theorem and I can't even get close to any of the options

OpenStudy (anonymous):

?

OpenStudy (anonymous):

http://imgur.com/Rz1JJCk

OpenStudy (anonymous):

the 2nd one? or the 1st one

OpenStudy (anonymous):

both? i want to learn how to do it

OpenStudy (anonymous):

ok let's see....

OpenStudy (anonymous):

the first one is wrong btw

OpenStudy (anonymous):

as u can see, the pathogrem threom is \(a^2+b^2=c^2\)

OpenStudy (anonymous):

oh yeah i randomly clicked that tbh i was getting frustrated

OpenStudy (anonymous):

whereas a and b is the height and the radius and c is the slant height

OpenStudy (anonymous):

so the first one... you plug the numbers in \(9^2+4^2=c^2\)

OpenStudy (anonymous):

i did 8^2+9^2 and got 145 and then i divided that in half so i got 72.5

OpenStudy (anonymous):

u will find it not a perfect square tho

OpenStudy (anonymous):

oh wait its 4 oh its half of the diameter ohhhhhhhhhhh

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

okay thank you so much ahh i'll try it now and see if i can get it

OpenStudy (anonymous):

ok.

OpenStudy (acxbox22):

not really

OpenStudy (anonymous):

yeah i know ^u^ i was about to edit that in

OpenStudy (anonymous):

lol

OpenStudy (acxbox22):

asking a good question does not involve attention grabbers

OpenStudy (anonymous):

:/ ok

OpenStudy (anonymous):

\(\tiny{I~suppose~so...}\)

OpenStudy (anonymous):

so I got 9.85, is that right?

OpenStudy (anonymous):

um hum

OpenStudy (anonymous):

or 9.8 bc im supposed to round it

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

now try the second one :P u can do this

OpenStudy (anonymous):

22.8?

OpenStudy (anonymous):

thats too high...

OpenStudy (anonymous):

um hum... remember.. the diameter...

OpenStudy (anonymous):

ohh right so 11.4

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

god ur amazing thank you :0

OpenStudy (anonymous):

ary i was trying to solve it using mental math

OpenStudy (anonymous):

its cool ^U^ mental math is hard when you have exponentss

OpenStudy (anonymous):

yeah but u don't really need to like solve solve it i know so it's cool ^o^

OpenStudy (anonymous):

any more?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

since when r u here sai?

OpenStudy (anonymous):

just now.. u have a problem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

good

OpenStudy (anonymous):

XD jk im not that mean

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

ah i think i have another one for u let me screenshot it

OpenStudy (anonymous):

http://i.imgur.com/3kURHRv.png?1

OpenStudy (anonymous):

oh sh*t i hate those problems but ok i'll try to help u

OpenStudy (anonymous):

it would be a lot easier if it were cubes man

OpenStudy (anonymous):

yeah same lmao

OpenStudy (anonymous):

i know i did something wrong bc i got 8.06

OpenStudy (anonymous):

find the diagonal of the base first

OpenStudy (anonymous):

tell me u know how to do that...

OpenStudy (anonymous):

lol ik how to do that!!!!!!!! b ut of course... im too lazy :P molly.. I BELIEVE IN U!

OpenStudy (anonymous):

so u like multiply the length of one side by the square root of 2 right

OpenStudy (anonymous):

i don't get u wht did u get as the answer tho?

OpenStudy (anonymous):

im lost tbh

OpenStudy (anonymous):

here hang on i think they gave us instructions

OpenStudy (anonymous):

... let's see... \(4^2+10^2=c^2\) solve this

OpenStudy (anonymous):

*sai askes: where did you get that from?*

OpenStudy (anonymous):

58?

OpenStudy (anonymous):

116=\(x^2\)

OpenStudy (anonymous):

http://imgur.com/5tg2i6Q

OpenStudy (anonymous):

116 / 2 = 58 = x right

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

the square root! not divided by 2!

OpenStudy (anonymous):

no u square root it not divide it by 2

OpenStudy (anonymous):

its \(\sqrt116\)

OpenStudy (anonymous):

sorry lmao my brain is being dumb i knew i was doing that wrong

OpenStudy (anonymous):

\(\sqrt{116}\)

OpenStudy (anonymous):

not 116 \(\div\) 2!!

OpenStudy (anonymous):

10.7

OpenStudy (anonymous):

jeez he got it

OpenStudy (anonymous):

so its about... 11

OpenStudy (anonymous):

round it!! don't make it harder!! make it easier so its 11

OpenStudy (anonymous):

ty u guys

OpenStudy (anonymous):

now solve this \(11^2+10^2=c^2\)

OpenStudy (anonymous):

well its multiple choice and im supposed to round it to the nearest tenth so 10.77 would be 10.8 there isnt even an option for 11

OpenStudy (anonymous):

221=\(c^2\)

OpenStudy (anonymous):

no it's not done yet... u found the diagonal of the BASE not the RECTANGULAR PRISM

OpenStudy (anonymous):

14.86 oh okay

OpenStudy (anonymous):

:Dnex t question

OpenStudy (anonymous):

=-= 15!

OpenStudy (anonymous):

the answer is 14.7 tho

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

u gtta ROUNDD welll..... true... but then.... D: I WANNA ROUND IT

OpenStudy (anonymous):

now for the enxt problem... solve this: \(14^2+3^2=c^2\)

OpenStudy (anonymous):

205=\(c^2\)

OpenStudy (anonymous):

somewhere around 14. so now solve this: \(14^2+4^2=c^2\)

OpenStudy (anonymous):

and then u can get ur answer

OpenStudy (anonymous):

14.31?

OpenStudy (anonymous):

i think it should be a bit more than that

OpenStudy (anonymous):

through mental math, i figured that it should be 14.8

OpenStudy (anonymous):

hat was it last time this answer is like 14.5

OpenStudy (anonymous):

14.56 so 14.7

OpenStudy (anonymous):

somethinglike that

OpenStudy (anonymous):

so 14.8? yeah immga go withh tha

OpenStudy (anonymous):

ugh lag